A 2400 baud, 7 bit word length, odd parity, and one stop bit message transmits characters per second. a) 100 b) 240 c) 440 d) 1000
Answer #1

Given:

2400 baud, 7-bit word length, odd parity, one stop bit

Answer:

Frame diagram with size:

Start Data Parity Stop
1 bit 5-9 bits 1 bit(either 0 or 1) 1-2 bits

2400: maximum of 2400 bits can be transferred per second

7-bit word length: data is being sent in blocks of size 7-bit

Odd parity takes a 1-bit size

Parity is set to 1 if the data bits contain an odd number of 1's, else it is set to 0

The stop bit is used to say that it is the end of the frame. It is of 1-bit in size.

total data to be sent = start bit + data bits + parity bit + stop bit = 1 + 7 + 1 + 1 = 10 bits.

Number of 7-bit word length data bits transmitted per second = 2400/10 = 240

Therefore, the correct option is b) 240